How do i solve, I know the answer is 8 but i dont know how to get it.
1/2+2/x=3/4

Answer by stanbon(75887)

1/2+2/x=3/4
-----
When you have denominators in an equation, get rid of them.
----
Multiply thru by 4x to get:
2x + 8 = 3x
=======
x = 8
==============
Cheers,
Stan H.
==============